OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 2/x+6 = 18/3x+18

OpenStudy (anonymous):

lests take x rational on one side

OpenStudy (anonymous):

4/3 = 2/x+6 - 18/3x+18

OpenStudy (anonymous):

now lets take rhs

OpenStudy (anonymous):

cross multiply it 2(3x+18)-18(x+6)/x+6(3x+18)

OpenStudy (anonymous):

so will be 6x+36-18x-108/3x^2+18x+18x+108

OpenStudy (anonymous):

therefore 4/3 = -12X-72/3X^2+36X+108

OpenStudy (anonymous):

Answer Choices: x = –11 x = –3 x = 3 x = 11

OpenStudy (anonymous):

now multiply 3x^2 eq. with 4 and 3 with -12x eq. 12x^2+144x+432 = -36x-216

OpenStudy (anonymous):

now its very easy there will be quad. eq. 12x^2+180x+648 now take out common to make it small x^2+15x+54

OpenStudy (anonymous):

now use quadratic formula or split method i prefer the formula so answer will be -3

OpenStudy (anonymous):

did u get it

OpenStudy (anonymous):

Yes thank youu